Suppose your waiting time for a bus in the morning is uniformly distributed on [0,

8], whereas waiting time in the evening is uniformly distributed on [0, 10] independent

of morning waiting time.


a. If you take the bus each morning and evening for a week, what is your total

expected waiting time? [Hint: Define rv's ?1, … , ?10 and use a rule of expected

value.]


b. What is the variance of your total waiting time?


c. What are the expected value and variance of the difference between morning

and evening waiting times on a given day?


d. What are the expected value and variance of the difference between total

morning waiting time and total evening waiting time for a particular week?
Answer #1

Answer

A bit theory

Solution

Here by the problem let us assume X_1,X_2,...,X_5 \ \ and \ \ Y_1,Y_2,...,Y_5 be the random variables representing waitimg times of the morning and evenings respectively where X_i\sim Uniform(0,8) \ \ ,i=1(1)5 \ \ and \ \ Y_j\sim Uniform(0,10) \ \ ,j=1(1)5 and all Xi's and Yj's are independent.

So then E(X_i)=(0+8)/2=4 \ \ and \ \ Var(X_i)=(8-0)^2/12=16/3

and ,Cov(X_i,X_{i'})=0 \ \ , for \ i\neq i' , due to independence

Similarly,

E(Y_j)=(0+10)/2=5 \ \ and \ \ Var(Y_j)=(10-0)^2/12=25/3

and ,Cov(Y_j,Y_{j'})=0 \ \ , for \ j\neq j' due to independence  

And on the other hand, Cov(X_i,Y_j)=0 for all i and j

(a) So the total waiting time be,

T=X_1+X_2+...+X_5+Y_1+Y_2+...+Y_5=\sum_{i=1}^5X_i+\sum_{j=1}^5Y_j

So the expected total waiting time be,

E(T)=E\left ( \sum_{i=1}^5X_i+\sum_{j=1}^5Y_j \right )=\sum_{i=1}^5E(X_i)+\sum_{j=1}^5E(Y_j)=5\times 4+ 5\times 5=45

(b) And the variance of total waiting time in 5 days be,

Var(T)=Var\left ( \sum_{i=1}^5X_i+\sum_{j=1}^5Y_j \right )=\sum_{i=1}^5Var(X_i)+\sum_{j=1}^5Var(Y_j)

since the mutual covariances are all zero due to independence.

=5\times (16/3)+5\times (25/3)=205/3\approx 68.33

hence the answer...............

Thank you............


answered by: ANURANJAN SARSAM